Use the Remainder Theorem to determine the remainder when P(x) is divided by D(x). Determine if D(x) is a factor of P(x).


  1. P(x) = x3 - x2 + x - 1; D(x) = x - 1
  2. P(x) = 20x3 - 17x2 + 6x - 26; D(x) = x + 1
  3. P(x) = 2x3 - 5x + x2 + 2; D(x) = x + 2
  4. P(x) = 12x3 - 8x2 + 4x - 3; D(x) = 2x - 1










1
Expert's answer
2021-01-28T05:08:55-0500

By The Remainder Theorem:

When we divide a polynomial "P(x)"  by "(x-c)" the remainder is "P(c)"

1. "P(x)=x^3-x^2+x-1;D(x)=x-1"


"P(1)=(1)^3-(1)^2+1-1=0"

"D(x)" is a factor of "P(x)."


2. "P(x)=20x^3-17x^2+6x-26;D(x)=x+1"


"P(-1)=20(-1)^3-17(-1)^2+6(-1)-26=-69"


Remainder is "-69."


3. "P(x)=2x^3+x^2-5x+2;D(x)=x+2"


"P(-2)=2(-2)^3+(-2)^2-5(-2)+2=0"

"D(x)" is a factor of "P(x)."


4. "P(x)=12x^3-8x^2+4x-3;D(x)=2x-1"


"P_1(x)=6x^3-4x^2+2x-1.5" divided by "D_1=(x-0.5)"



"P_1(0.5)=6(0.5)^3-4(0.5)^2+2(0.5)-1.5=-0.75"


Remainder is "2(-0.75)=-1.5."